the length of a rectangle is one more than double the width. If the perimeter is 32 kilometers, what are the length and width?
Answer by ReadingBoosters(3246)

l = 2w+1
p=2(l+w) = 32
2((2w+1)+w)=32
4w+2+2w=32
6w+2=32
6w=30
w=5 km
l=2(5)+1=11 km
..........
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.